A store sells bluegrass seed for $6 per pound and ryegrass seed for $3 per pound. How much ryegrass must be mixed with 90 pounds of bluegrass to obtain a blend that will sell for $5 per pound?
Answer by richwmiller(17219)

3*x+6*90=5(90+x)
3*x+540=450+0.05x
3x-5x=450-540
-2x=-90
x=45 lbs of $3 is added to 90 lbs of $6 to get 135 lbs at $5
check
3*45+6*90=5(90+45)
135+540=5(135)
675=675
ok